OpenStudy (lena772):

A computer manufacturer built a new facility for assembling computers. There were construction and new equipment costs. The company paid for these costs and made combined profits of $40 million after 4 years while profits increased $30 million per year. Select the correct graph of this function. There is a graph of a line that has a y intercept that is approximately $80 million and passes through the point of approximately x = 2 years and y = $140 million.

OpenStudy (lena772):

There is a graph of a line that has a y intercept that is approximately negative $80 million and an x intercept that is approximately 2.7 years. There is a graph of a line that has a y intercept that is approximately negative $45 million and an x intercept that is approximately 2 years. There is a graph of a line that has a y intercept that is approximately negative $45 million and an x intercept that is approximately 0.6 years.

ganeshie8 (ganeshie8):

looks u did this before :o

OpenStudy (lena772):

it's a little different

ganeshie8 (ganeshie8):

Yeah i see

OpenStudy (lena772):

how did you find the two points in the last one?

ganeshie8 (ganeshie8):

we dint find two points in last one

ganeshie8 (ganeshie8):

we're given a point and slope

ganeshie8 (ganeshie8):

read the problem and make a guess wats the point ? whats the slope ?

OpenStudy (lena772):

point (4, 40)

OpenStudy (lena772):

slope 30?

ganeshie8 (ganeshie8):

Correct ! write the equation and find x,y intercepts

OpenStudy (lena772):

y=30x+b 40=30*4+b 40=120+b -80=b

OpenStudy (lena772):

y=30x-80

OpenStudy (lena772):

0=30x-80 80=30x 2.7=x

OpenStudy (lena772):

so b?

ganeshie8 (ganeshie8):

Yes ! good job !!